OpenStudy (anonymous):

13x+45y=4x+13y prove RHS=LHS

OpenStudy (anonymous):

\[13x + 45 y = 4x + 13y\] \[9x = -32 y\] x = -32 y /9 \[13\frac{ -32y }{ 9} + 45y = 4\frac{ -32y }{ 9} + 13 y\] \[= \frac{ -416 y}{ 9 } + 45 y = -128y/9 + 13y\] \[=> 32y = \frac{ 416y }{ 9 } -\frac{ 128y }{ 9 }\] \[=>32y = \frac{ 288y }{ 9 }\] \[32y = 32y\] Hence,It is proved that LHS = RHS @suzzmteazz

OpenStudy (anonymous):

@suzznteazz